The electric filed due in a static electric charge is 3 000 N/C and passes through a surface area of 0.3 m2. Calculate the electric flux

1
Expert's answer
2022-04-11T12:02:44-0400

The electric flux

"\Phi=\oint EdA=\frac{q}{\epsilon_0}"

"\Phi=\frac{3000}{8.85*10^{-12}}=3.4*10^{14}\:\rm N\cdot m^2/C"